# Scaling Reproducibility: An AI-Assisted Workflow

### v1

@@ 2.3 Reproducibility Criterion and Evaluation
For the original 67-study corpus, exact agreement with the manually replicated 2SLS coefficient serves as the benchmark for reproducibility. In other words, for these 70 specifications, the equality between the published coefficient and the manually replicated coefficient (up to rounding error) has already been established. Our objective is therefore to verify the second equality: that the harmonized pipeline reproduces the same 2SLS estimate.
For the newly incorporated studies, reproducibility has not been manually verified. Although the journals from which these studies are drawn typically require exact numerical replication as a condition of publication, we do not independently verify the equality between the published and authors’ pipeline outputs. Instead, we focus on achieving and documenting the second equality under our workflow. Verifying the first equality would sometimes require access to detailed Supplementary Materials, which are not always available, and is left for future work. Note that, as in Lal et al. (2024), reproducibility does not equal robustness or credibility, which requires additional diagnostics and knowledge about the research design.
Focusing on the 2SLS point estimate isolates the core execution bottleneck in large-scale reproducibility. Successful reproduction implies that the workflow has correctly identified the instrument set, endogenous regressor, control variables, fixed effects, sample restrictions, weights, and data transformations required for the model. Because the 2SLS coefficient is jointly determined by these components, exact numerical agreement indicates that the specification, data processing, and execution have been correctly harmonized. Conditional on this success, downstream diagnostic statistics are mechanically determined by the template. Reproducing the 2SLS point estimate therefore marks the critical breakpoint.
Having defined this criterion, we evaluate performance at both the paper and specification levels. We report ingestion success rates, specification extraction rates for up to three IV models per study, code execution rates, exact 2SLS replication rates at the specification level, and diagnostic report generation rates.
Because the pipeline is adaptive and version-controlled (see below), recurring execution failures are encoded as generalized rules between runs. As a result, conditional on accessible replication materials, stage-level success converges to full coverage under a fixed pipeline version. The metrics therefore assess whether the finalized executor can reliably harmonize heterogeneous replication packages and scale the 2SLS–OLS diagnostic analysis both forward in time and within studies.
3. AI-Assisted Reproducibility Workflow
This section describes the AI-assisted workflow used to execute the reanalyses of IV corpus defined in Section 2. The workflow targets a practical bottleneck in reproducibility: executing established research protocols reliably across heterogeneous replication packages while preserving numerical precision and auditability. This workflow does not automate methodological reasoning or introduce new statistical procedures. Instead, it standardizes and accelerates execution when usable data and code are available.
3.1 Design Principles
Large-scale reproducibility involves a basic tension between heterogeneity and determinacy. Replication materials vary widely across studies. Even within the IV corpus, papers differ in programming language (Stata, R, and Python), directory structure, naming conventions, and documentation quality. At the same time, reproducibility requires determinacy: for a fixed pipeline version and fixed inputs, numerical outputs must not depend on ad hoc decisions, platform-specific defaults, or stochastic behavior.
The workflow resolves this tension by separating adaptive coordination from fixed computation. Adaptation is used to route tasks, interpret failures, and select among predefined recovery steps. Numerical work—data preparation, model estimation, and diagnostic computation—is executed by version-controlled program code. For a fixed pipeline version and fixed inputs, the workflow produces identical numerical outputs and retains a complete audit trail of intermediate artifacts and logs. When new failure patterns are encountered, fixes are incorporated between runs and version-controlled, rather than allowing numerical behavior to drift within runs. Note that uncertainty estimates, particularly those based on bootstrap or jackknife procedures, make exact replication more challenging. This is because random seed behavior is not always portable across platforms, and parallel computation can further complicate seed control.
A second principle is the separation of scientific reasoning from execution. As described in Section 2, the diagnostic template—based on Lal et al. (2024) and ivDiag—is fixed in advance. Once this template is specified, reproduction reduces to execution-oriented tasks: acquiring replication packages, reconstructing computational environments, identifying IV specifications, running code, extracting estimates, and compiling diagnostic outputs. The workflow evaluates whether these tasks can be automated without loss of precision.
3.2 Architecture and Implementation
We implement the workflow using Claude Code Skills (hereafter, Skills), an agentic system organized as a three-layer architecture. At the top layer, an LLM (Claude) serves as an orchestrator that dispatches tasks, interprets errors, and determines how the pipeline proceeds. The middle layer consists of structured skill descriptions that define each stage’s input–output contract and record previously resolved failure patterns. At the bottom layer, rule-based agent code and diagnostic scripts perform all file operations and statistical computation. The upper layers govern coordination and adaptation; the bottom layer governs numerical results.
The orchestrator reads task instructions, consults the relevant skill descriptions, writes and edits plain-text artifacts (primarily Markdown configuration files and small Python utilities), and invokes external tools as needed. It does not perform statistical estimation. All estimation and diagnostics are executed by explicit code in R, Stata, and Python. In particular, the full diagnostic suite is implemented in a standalone R script that operates solely on exported analysis datasets. The LLM therefore controls task routing and structured interpretation, but it is excluded from numerical estimation and inference. Figure 1 illustrate the three-layer architecture.
Reproduction is decomposed into modular stages because replication packages vary widely in software, directory structure, naming conventions, and documentation. Failures typically arise at distinct points—retrieval, parsing, code repair, execution, or reporting—and require stage-specific information. Each stage is assigned to a dedicated agent. Agents communicate exclusively through standardized intermediate files written to disk (e.g., JSON, CSV, logs) and share no hidden state. Each stage reads explicit inputs and writes explicit outputs, making every step inspectable and rerunnable. When a failure occurs, execution can resume from the affected stage without restarting the entire pipeline.
The system was developed iteratively through repeated encounters with diverse replication packages. When a new failure pattern is identified, the orchestrator proposes a diagnosis and candidate fix. After review, successful fixes are incorporated as version-controlled updates to the execution layer and implemented between runs rather than during execution. Each run is therefore tied to a fixed, inspectable pipeline version, and expanded coverage does not alter the numerical behavior of prior versions.
Figure 1: Overview of the agentic AI workflow for reproducibility. The above figure illustrates the three-layer agentic architecture enabled by Skills. A top-layer LLM orchestrator routes tasks and interprets errors but does not perform estimation. The middle layer defines structured input–output contracts and records resolved failure patterns. The bottom layer consists of rule-based agent code and diagnostic scripts in R, Stata, and Python that execute all file and statistical operations through a modular seven-stage pipeline, from material acquisition to standardized reports.
The pipeline comprises seven stages, also illustrated in Figure 1. Given a paper in PDF format, an agent called Profiler extracts metadata and replication links. The Librarian downloads associated data and code. Then, the same Profiler agent identifies IV specifications within the replication code, including up to three targeted specifications per study as defined in Section 2. The Janitor prepares scripts for execution by resolving path dependencies and environment assumptions. The Runner executes the models and extracts 2SLS estimates. The Skeptic applies the fixed diagnostic template for IV designs. The Journalist compiles standardized reports. Each stage writes inputs and outputs to disk and records runtime, status, and error messages in structured logs, which form the basis for the stage-level performance metrics reported in Section 5.
3.3 Challenges in a Skills-Based Workflow
Using a Skills-based orchestration layer introduces practical challenges that are easy to understate if one only describes the workflow at a high level. We highlight four challenges that arise specifically because task routing and failure handling are AI-assisted, and we summarize how our implementation constrains these risks.
First, agent-based workflows can introduce run-to-run variation through branching logic. Even when numerical code is fixed, differences in routing or recovery steps can change which scripts are executed, which intermediate datasets are exported, or which specification is treated as primary. We constrain routing decisions to explicit, inspectable artifacts and fixed rules wherever possible. Agents define subsequent inputs through standardized files, and the orchestrator selects among documented options while recording all decisions in logs. For a fixed pipeline version and fixed inputs, reruns produce identical numerical outputs.
Second, coverage can expand in ways that are difficult to audit. Without discipline, a workflow may accumulate undocumented heuristics or paper-specific patches. We address this by version-controlling all deterministic code and recurring resolution rules. When a new failure pattern is identified, the fix is implemented as a generalized update between runs rather than as an ad hoc intervention within a single execution. Each run is therefore anchored to a stable pipeline state.
Third, execution variability across replication packages generates complex failure modes. Differences in software environments, directory structures, naming conventions, data encodings, and multi-language codebases create errors that cannot be fully anticipated ex ante. Some failures do not produce explicit error messages but instead appear as inconsistencies across intermediate artifacts, such as mismatched coefficients or incomplete exports. The workflow addresses these patterns through a structured resolution cycle. The orchestrator inspects logs and cross-validates outputs, traces failures across stages to identify root causes, and implements fixes as generalized rules in deterministic code. Resolved patterns are recorded in a structured knowledge base, allowing solutions to persist across runs while preserving numerical determinacy within each version.
Finally, supporting additional research designs creates pressure to embed design-specific assumptions in the executor. We avoid this by keeping the executor design-agnostic and treating diagnostic templates as inputs. Researchers specify estimands, estimators, and summary measures in the template. The executor performs the same execution tasks—acquisition, preparation, execution, extraction, and reporting—while invoking the appropriate template. Extending to new designs therefore requires modifying templates rather than rewriting the underlying architecture.
4. Demonstration
The previous section described each pipeline stage in abstract terms. Here we trace a single AJPS study, Rueda (2017), through the full workflow, from the input PDF to the final diagnostic report. At each stage, we show the intermediate artifacts the system generates, illustrating how the modular, file-based architecture supports auditability.
This demonstration reflects the cumulative knowledge embedded through the adaptive mechanism described above. To name a few nontrivial challenges the workflow has learned to handle: parsing the ivreg2 syntax in the author’s Stata code; resolving time-series operators such as l. and l4. embedded in variable names; enforcing the e(sample) post-estimation restriction; and executing a 13-script build chain of .do files. These capabilities were acquired from recurring patterns in earlier papers, indicating that the accumulated knowledge generalizes beyond the initial development set.
Rueda (2017) studies the relationship between polling station size and vote buying in Colombia, instrumenting average polling station size with the maximum size set by election authorities. This case is well suited for demonstration for three reasons. First, the design is straightforward, with one endogenous regressor and a clearly defined instrument. Second, the paper reports three IV specifications with strong first stages and adequate power, consistent with the diagnostics in our template. Third, the replication code is written in Stata, the dominant environment in our evaluation corpus and comparatively difficult for AI systems to parse and execute due to its closed-source ecosystem and idiosyncratic syntax. We now walk through the pipeline step by step.
Stages 1–2: Acquire Materials
The workflow begins with the paper’s PDF as the only input. In this first stage, the Profiler parses the PDF and extracts structured metadata, including the title, author, journal, and, critically, the data repository URL embedded in the data availability statement. The output is a JSON file:
1ivreg2 e_vote_buying l4.margin_index2 l.nbi_i ///
2 l.own_resources lpopulation l.armed_actor ///
3 l4.lsize lpotencial ///
4 (lm_pob_mesa = lz_pob_mesa_f) if e(sample), ///
5 first cluster(muni_code)
In Stage 2, the Librarian resolves the Dataverse DOI, queries the Dataverse API for the full file listing, and downloads the complete replication package. The materials include multiple Stata data files (.dta), dataset-construction scripts, analysis scripts, and a master analysis file. All files are stored in a versioned workspace directory. The download log confirms that both the article and the replication package were retrieved without error.
Stage 3: Determine Specifications
With the replication materials in place, the Profiler parses all Stata scripts to identify IV estimation commands. In the main analysis file (main_results_aggregate_data.do), it detects three ivreg2 calls—a user-contributed two-stage least squares routine with syntax distinct from the built-in ivregress 2sls. Support for ivreg2 was added incrementally as similar patterns were encountered in earlier papers. The first command appears as follows:
1{
2 "title": "Small Aggregates, Big Manipulation: Vote Buying Enforcement
3 and Collective Monitoring",
4 "authors": "Miguel R. Rueda",
5 "year": "2017",
6 "journal": "American Journal Of Political Science",
7 "replication_url": "http://dx.doi.org/10.7910/DVN/K6ZOOW"
8}
A central feature of this specification is the use of Stata time-series operators. For example, l.nbi_i denotes the first lag of nbi_i, and l4.margin_index2 denotes the fourth lag. These operators must be mapped to realized variable names in the constructed dataset. The Profiler and Runner resolve this mapping using translation rules accumulated during prior executions.
For each ivreg2 call, the Profiler constructs a structured representation that records the outcome, endogenous regressor, instrument, controls, sample restriction, and clustering variable. In this case, all three specifications share the same treatment and instrument, cluster at the municipality level, and differ only in the outcome variable or the set of controls. Rather than reproducing the full table of variables, we emphasize that the extracted structure captures only role assignments and estimation options necessary for downstream execution and diagnostics.
All extracted fields are written to metadata.json, which serves as the contract between the identification stage and subsequent execution and evaluation stages.
Stages 4–5: Code Preparation and Execution
The original replication package is organized as a multi-file Stata project. Dataset-construction scripts generate intermediate .dta files, which are subsequently loaded by analysis scripts to produce the reported tables. The IV specifications are embedded within this build chain rather than implemented in a single script. Reproducing the results therefore requires executing the scripts in the correct order and preserving dependencies across files.
Before execution, the workflow harmonizes the code for automated use. The Janitor redirects file paths to the local workspace and removes interactive or side-effect commands such as graph export and log using, which would otherwise interrupt batch execution. It also instruments each ivreg2 call so that coefficients and standard errors can be programmatically extracted from the log. These adjustments do not alter the statistical content of the original code; they standardize the environment for reproducible execution. All modifications are recorded in structured JSON logs, preserving a complete audit trail of the harmonization process.
A substantive complication arises from the use of if e(sample) in the original Stata commands. This condition restricts the estimation sample to observations selected in a prior regression. Preserving the author’s intended sample requires respecting this sequencing: the conditioning regression must be executed first, and data export must occur before subsequent commands overwrite the sample indicator. Correct handling of this dependency is essential for reproducing the reported specifications.
The Runner then executes the harmonized build chain in batch mode. All scripts complete without error. The system extracts the IV estimates from the marked log output and exports the corresponding analysis datasets for downstream diagnostics. As a cross-language validation step, each specification is re-estimated in R using iv_robust() from the estimatr package. The resulting coefficients match the Stata outputs within numerical tolerance, confirming that the harmonized execution preserves the original estimates.
Stage 6: Diagnostic Analysis
The Skeptic reads metadata.json to recover the model specification and loads the corresponding analysis_data_spec_*.csv files. After matching variable names to dataset columns, it invokes the R diagnostic script (diagnostics_core.R), which implements the full set of procedures described in Section B in the Supplementary Materials. Because the original IV designs use a single instrument, the script also applies the tFtF procedure (Lee et al., 2022), which adjusts critical values for the conventional tt-test as a function of the first-stage FF-statistic. Table 2 reports the results.
Table 2: Diagnostic results for Rueda (2017)
Specificaiton 1 Specificaiton 2 Specificaiton 3
Outcome variable e_vote_buying sum_vb e_vote_buying
Treatment variable lm_pob_mesa lm_pob_mesa lm_pob_mesa
Instrument lz_pob_mesa_f lz_pob_mesa_f lz_pob_mesa_f
Clustering varaible munni_code munni_code munni_code
#Covariates 7 7 2
Instrument strength
     Effective FF 827.2 203.9 8,598.3
     Bootstrap FF 925.5 202.6 8,989.5
2SLS estimate
     Coefficient -1.460 -2.242 -0.984
     Std. error 0.463 1.300 0.142
     pp-value for tt-stat 0.002 0.085 0.000
     #Observations 4,352 1,069 4,352
     #Clusters 1,098 632 1,098
Robust inference
     AR pp-value 0.002 0.075 0.000
     tFtF p<0.05p<0.05 Yes No Yes
     Bootstrap-cc 95% CI [-2.42, -0.62] [-4.66, -0.07] [-1.32, -0.73]
     Bootstrap-tt 95% CI [-2.37, -0.55] [-5.74, 1.26] [-1.23, -0.73]
     Any CI includes 0? No Yes: Boot-tt No
Sensitivity (jackknife)
     Range [-1.49, -1.31] [-2.27, -0.94] [-0.99, -0.96]
@@ 4. Demonstration
OLS comparison
     OLS coefficient -0.626 -0.984 -0.675
     2SLS/OLS ratio 2.3 2.3 1.5
Robustness rating HIGH MODERATE HIGH
Notes: All bootstrap tests use 1,000 iterations with cluster-level resampling. Jackknife analysis removes one municipality cluster at a time. The effective FF-statistic (Montiel Olea and Pflueger, 2013) is reported as the primary measure of instrument strength. Spec 2 triggers two warnings: (1) the Anderson–Rubin pp-value exceeds 0.05; (2) removing municipality 11001 (Bogotá) changes the estimate by 58%, exceeding the 20% sensitivity threshold.
All three specifications display strong first stages. The effective FF-statistics range from 204 to 8,598, well above the conventional cutoff of 10, and no model is flagged for weak instruments. Maximum polling station size is therefore a strong predictor of average station size in each specification.
For Specifications 1 and 3, the 2SLS estimates are statistically significant using the conventional tt-test, bootstrap methods, tFtF test, and Anderson-Rubin test, and the jackknife ranges are tight. The estimates from the full and reduced control sets are similar in sign and magnitude, providing an internal robustness comparison within the paper’s design.
For Specification 2, the instrument strength remains high, but inference is less robust. The pp-value for the Anderson-Rubin test exceeds 0.05, the tFtF test does not reject, and the bootstrap-tt interval includes zero. The jackknife identifies municipality 11001 as influential; removing it shifts the estimate by 58%. Under the prespecified rules, these two warnings yield a MODERATE rating. The system does not offer a substantive interpretation. It applies the same criteria to each specification and reports the results. Whether the weaker evidence reflects lower power in the smaller sample or a different relationship for the alternative outcome remains a question for the researcher.
Note that the reanalysis does not assess the credibility of the core identification assumptions, namely unconfoundedness and the exclusion restriction of the instrument. When there is a single instrument for a single endogenous regressor, these assumptions are not directly testable, and the diagnostics cannot adjudicate their validity.
Stage 7: Report Findings
The Journalist then assembles the outputs from Stage 6 into a standardized report, report.pdf, which is included in the Supplementary Materials (Section E). The report is organized by specification. It begins with an executive summary table that lists the trustworthiness ratings. It then documents the study design, variable definitions, replicated IV estimates, and the full set of diagnostics. For each specification, four figures are generated automatically: following Lal et al. (2024), a coefficient comparison plot displaying OLS and 2SLS estimates with multiple confidence intervals, a comparison of first-stage FF-statistics, bootstrap confidence intervals, and jackknife sensitivity. The format is identical across papers, which facilitates cross-study comparison.
(a) Executive summary
(b) Coefficient plot for Spec 1
Figure 2: Selected figures from the diagnostic report for Rueda (2017) generated by the AI workflow. Left: the executive summary page summarizing ratings for all three specifications (two HIGH, one MODERATE). Right: the coefficient comparison plot for specification 1, showing OLS and 2SLS point estimates with analytic, bootstrap-cc, bootstrap-tt, tFtF, and Anderson-Rubin confidence intervals.
The full pipeline—from PDF ingestion and downloading the replication package to report generation—completed in less than four minutes. Diagnostic computation in Stage 6 accounts for most of the runtime. No human intervention was involved. Click [here] for a real-time demonstration, although the implementation itself does not rely on a graphical interface.
5. Main Findings
This section reports the main findings from implementing the AI-assisted replication workflow at scale. We first evaluate the workflow’s performance, including its success rate, binding constraints, and execution time. We then present the empirical results from the extended IV corpus, showing that the automated pipeline reproduces the core patterns documented in Lal et al. (2024) under the same diagnostic template.
5.1 Performance of the AI Workflow
We evaluate the workflow on the 67 IV papers that form the benchmark corpus in Lal et al. (2024) and on 25 newly collected studies under identical inclusion criteria. Table 3 reports stage-level and end-to-end success rates at both the paper and specification levels.
For the original sample, end-to-end autonomous success is 55/67 (82%). All failures occur at the material retrieval stage, where replication archives are no longer publicly available. These 12 papers—all published before 2020—cannot currently be downloaded from public repositories. Using archived materials retained from the earlier project, however, we execute the pipeline and reproduce their findings. Conditional on accessible data and code, specification extraction, execution, and diagnostic analysis succeed for all papers. Thus, among studies with available materials, the current pipeline achieves full coverage.
Table 3: Success Replication Rate (Original and Expanded Samples)

### v2

@@ A.1 IPW Estimator and Variance Formula
[…]
Under stratified simple random sampling without replacement within each cell, the variance is
[…]
Table S1: Estimated Reproducibility Rates: Cluster Bootstrap Confidence Intervals
Journal All Years Pre-DA-RT Post-DA-RT
APSR ($\geq$2016) Rate 61.2% 16.9% 73.0%
95% CI [46.9, 70.9] [12.5, 21.2] [64.9, 79.1]
Papers sampled 128 48 80
AJPS ($\geq$2015) Rate 71.8% 47.0% 84.4%
95% CI [59.3, 82.9] [28.0, 62.5] [76.5, 92.2]
Papers sampled 128 40 88
JOP ($\geq$2015) Rate 67.8% 20.6% 81.2%
95% CI [52.6, 80.4] [5.1, 35.1] [71.3, 91.3]
Papers sampled 128 40 88
Pooled Rate 67.2% 29.6% 79.8%
95% CI [58.8, 74.4] [18.0, 40.9] [74.4, 85.4]
Papers sampled 384 128 256
Notes. Point estimates are identical to Table 4. Confidence intervals use the percentile method from 5,000 cluster bootstrap resamples of journal-year cells (seed 20260322). “Pre-DA-RT” and “Post-DA-RT” use journal-specific cutoffs: APSR $\geq$2016, AJPS $\geq$2015, JOP $\geq$2015. Bootstrap intervals are wider for individual journals (5–11 clusters per period) but closely match the analytic intervals for pooled estimates.
A.2 Diagnostic Classification
Tables LABEL:tab:diagnostic-gap and LABEL:tab:diagnostic-repaired provide a paper-level classification of all 25 papers that received a verdict below fully reproducible at any point during processing, as summarized in Section 5.2. Table LABEL:tab:diagnostic-gap lists the 15 papers whose gaps remain unresolved—due to missing author data, pipeline limitations, or ambiguous code-to-table mappings. Table LABEL:tab:diagnostic-repaired lists 10 papers where the workflow’s adaptive execution repaired author-side packaging errors (missing dependencies, deprecated syntax, broken paths), recovering them to fully reproducible or near-fully reproducible status.
Table S2: Papers Remaining Below Fully Reproducible
Paper Verdict Category Why not fully reproducible
Author-caused unresolved gaps
Paper 1 Partial Missing author data A main-text table requires a data file not included in the released package; only appendix or intermediate files are provided.
Paper 2 Partial Restricted author data The README states that the full analysis depends on restricted registry data; only a subset of results is reproducible from released summary statistics.
Paper 3 Partial Missing data object A main-text table depends on a data object not shipped in the public package; the released file contains only pre-computed summaries.
Paper 4 Partial Missing code/spec. The released package exposes a richer model specification but not the code needed to reproduce a reduced model in the main table.
Paper 5 Partial Missing author data The package omits a survey source file required for one of the main-text tables.
Paper 6 Partial Paper/code divergence The shipped code runs and produces model output, but the remaining gap is between paper-reported values and the code-generated coefficients.
Paper 7 Not repl. Missing data inputs One set of main-text tables replicates, but remaining tables require cleaned input files absent from the released package.
Paper 8 Restricted Proprietary data Key variables were purchased from a commercial provider and excluded from the public release, preventing construction of the central exposure measures.
Paper 9 Restricted Permissioned data The replication package requires variables from a government agency that are not included in the release.
Paper 10 Largely Corrupt data file The main-text do-file aborts because a required data file is unreadable, blocking remaining table columns before estimation begins.
Pipeline-side deep semantic cases
Paper 11 Not repl. Deep semantic recon. Raw models run and some transformed quantities are visible, but most remaining paper rows depend on ambiguous post-estimation transforms that cannot be exposed safely without deeper reconstruction.
Paper 12 Partial Rendered-table recon. The missing table columns come from rendered output; recovered model families are not on the same displayed scale as the paper columns.
Paper 13 Partial Incomplete rendered recovery Printed regression-table output recovers some missing columns but does not provide a safe one-to-one reconstruction of the remaining table family.
Paper 14 Largely Matching-pipeline recon. The missing matched-model columns require a saved matched dataset and weighting artifact that are absent; the script does not complete the computationally intensive matching stage in batch.
Ambiguous cases
Paper 15 Not repl. Appendix vs. main-table mapping The available code prints an appendix matrix rather than the paper’s main table, and there is no safe local proof that the appendix outputs correspond to the unmatched main-text rows.
Table S3: Author-Side Packaging Errors Repaired by the Workflow
Paper Verdict Category What was repaired
Paper 16 Fully Missing dependencies The package assumed user-written Stata dependencies without a working local setup; installing them recovered the core tables.
Paper 17 Fully Fragile output layer Rendering commands stopped execution before results were fully exposed; bypassing those blockers recovered the paper.
Paper 18 Fully Missing Stata dependency The released script called a user-written command without shipping or installing it; isolating the failing script and running the remaining regressions recovered the paper.
Paper 19 Fully Missing Stata dependencies Legacy user-written commands were required for the last two main-text models; locating and installing them recovered the missing table columns.
Paper 20 Fully Missing data & deprecated output The package referenced a missing data file and relied on deprecated output assumptions; isolating the broken script recovered the viable analysis chain.
Paper 21 Fully Deprecated syntax Deprecated loop constructs and brittle setup paths blocked the main analysis pipeline; repairing those issues recovered the paper.
Paper 22 Fully Missing Stata dependency The script depended on a user-written package for auxiliary output; bypassing the failing script preserved the substantive regression pipeline.
Paper 23 Fully Missing sub-script The master do-file called a sub-script absent from the released package; running the surviving scripts directly recovered all main-text outputs.
Paper 24 Fully Column-name bug The released analysis script expected a column name not present in the shipped data; correcting the mismatch recovered the main table models.
Paper 25 Partial Fragile RD bookkeeping Legacy post-estimation bookkeeping after valid RD estimates crashed the loop; stabilizing those blocks allowed most substantive models to run.
A.3 IV Pipeline Performance
Table S4 reports stage-level and end-to-end success rates for the IV diagnostic corpus described in Section 5.3. Because failures are cumulative, the table reports both paper-level and specification-level outcomes.
Table S4: IV Corpus: Stage-Level Success Rates